The demand for apple orchard workers is given by the equation W = 80 - 3 Q and the supply



is given by W = 30 + 2 Q where W is the daily wage rate and Q is the quantity of workers.



Calculate Elasticity of demand and supply at equilibrium price and quantity.

1
Expert's answer
2022-02-28T14:01:39-0500

"W=80-3Q"

"W=30+2Q"

At equilibrium, demand=supply;

so, "80-3Q=30+2Q"

like terms, 5Q=50;

"\frac{\cancel5Q}{\cancel5}=\frac{50}{5}"


"Q=\frac{50}{5}=10"


We can find W by substituting Q with 10 in one of the equations.

"W=80-(3\times10)=50"


So, equilibrium point is at 50 for W and 10 for Q.


Elasticity:

"Change in W=50-0=50"


"Average of W=\frac{50+0}{2}=25"


"Percentage change=\frac{50}{25}\times100=200."

Change in Q"Change in Q=10-0=10"


"Average of Q=\frac{10+0}{2}=5"


"Percentage change=\frac{10}{5}\times100=200."


"Elasticity=\frac{percentage change in W}{percentage change in Q}=\frac{200}{200}=1"


Elasticity is 1.